# Beyond Early-Token Bias: Model-Specific and Language-Specific Position Effects in Multilingual LLMs

### v2

@@ Abstract
Large Language Models (LLMs) exhibit position bias—a systematic tendency to neglect information at specific context positions. However, the patterns of position bias behavior, depending on the language or model, remain unexplored. We present a multilingual study across five typologically distinct languages (English, Russian, German, Hindi, and Vietnamese) and five model architectures, examining how position bias interacts with prompt strategies and affects output entropy. Our key findings are: (1) Position bias is primarily model-driven, yet exhibits language-specific variations. For instance, Qwen2.5-7B-Instruct and DeepSeek 7B Chat consistently favors late positions, challenging established assumptions of a universal early-token bias in LLMs. (2) Explicitly instructing the model that "the context is relevant to the query" unexpectedly reduces accuracy across languages, undermining common prompt-engineering practices. (3) While the largest accuracy drop occurs when relevant information is placed in the middle of the context, this is not explicitly reflected by a corresponding peak in output entropy.
@@ 1 Introduction
Many recent applications based on large language models (LLMs) require support for long-context processing [^1] Such applications may include Retrieval-Augmented Generation, autonomous agents, scientific research, customer support, among others.. However, developing new training strategies to accommodate longer contexts is insufficient, as new challenges continue to arise. One notable issue is position bias – the systematic neglect of information located at specific positions, typically in the middle of the context (Baker et al., 2024)
While position bias has been well-documented in English-centric studies (Zhang et al., 2024b; Baker et al., 2024), its manifestation in multilingual contexts and across-various architectures remains underexplored. Furthermore, existing bias mitigation strategies (Peysakhovich & Lerer, 2023; Zhao et al., 2021; Zhang et al., 2024a; Guo & Vosoughi, 2024) have predominantly been evaluated on English-language datasets. As discussed in Section 2, lexical and morphological variations across languages necessitate careful consideration in multilingual modeling. As a step toward adapting multilingual large language models for real-world long-context applications, this work seeks to address the following research questions:
Q1: Is position bias primarily a model-driven phenomenon, or do language-specific patterns emerge due to lexical, morphological, and syntactic differences (Ghosh et al., 2024b)?
Q2: Do prompt-based strategies (e.g., explicit position guidance (Zhang et al., 2024a)) effectively mitigate bias across languages? Can we focus model attention via prompting, and would it improve model performance?
@@
Contributions. We present a multilingual analysis of position bias across five typologically diverse languages (English, German, Russian, Hindi, and Vietnamese) and five model architectures. Our key contributions are:
  • We show that position bias is predominantly model-driven, yet exhibits significant language-specific variations. For instance, models like Qwen2.5-7B-Instruct (Qwen et al., 2025) and DeepSeek-7B-Chat (DeepSeek-AI, 2024) display a strong late-position bias, contradicting prior claims of an inherent early-token preference in LLMs (Wu et al., 2025; Barbero et al., 2025). In contrast, Llama3.1-8B-Instruct (Touvron et al., 2023) prioritizes early positions. We speculate these differences stem from variations in training data and model architecture.
  • We demonstrate that explicitly instructing models about correct context placement (e.g., “the correct context has label 1”) consistently degrades accuracy across all languages. This finding challenges established prompt-engineering practices (Zhang et al., 2024a).
@@
    Retrieval-augmented generation (RAG) systems must account for specific model and language characteristics. Strategies that reorder context based on the assumption that models prioritize recent or initial tokens (see Section 2) may be ineffective or detrimental for models with a late-position bias.
@@
    The observed entropy dynamics complicate uncertainty-based bias mitigation techniques (Duan et al., 2024) and are crucial for developing effective uncertainty quantification strategies in RAG pipelines.
@@ 2 Related Work
What causes position bias? Prior work identifies multiple contributing factors. (Zhang et al., 2024b) attribute position bias to U-shaped attention patterns in transformers, which prioritize extremal positions and degrade performance for mid-context evidence. Theoretical and empirical studies further demonstrate that transformers attention is inherently biased toward earlier tokens (Wu et al., 2025; Barbero et al., 2025). (Wu et al., 2025) explains position bias due to "causal masking inherently biases attention toward earlier positions, as tokens in deeper layers attend to increasingly contextualized representations of earlier tokens". Our results reveal exceptions where later positions are desired, highlighting the complexity of the problem. Training data biases, such as serial-position effects in corpora, shape how models prioritize sequence positions (Wu et al., 2025; Guo & Vosoughi, 2024).
@@
Interplay of Culture, Language, and Model Design. The way we perceive the world is influenced not only by our culture (Oyserman & Lee, 2008) but also by the language we speak (Boroditsky et al., 2003). The latter point is particularly relevant for LLMs, since they are trained on specific languages. Recent studies have shown that multilingual LLMs often initiate their "thinking process" in English, pivoting to the prompt’s original language in the middle layers (Zhang et al., 2024b; Peysakhovich & Lerer, 2023). These models exhibit lower lexical naturalness in non-English languages, with the naturalness gap being more pronounced for languages structurally distant from English (Guo et al., 2024). While the volume of training data plays a crucial role (Arnett & Bergen, 2025), linguistic complexity — including lexical and morphological variations across languages — must also be considered (Ghosh et al., 2024b; Dang et al., 2024; Ismayilzada et al., 2025). Additionally, architectural design choices affect languages in different ways; for instance, removing positional encoding from language models would most degrade performance in languages with limited morphological systems (Ghosh et al., 2024b).
@@
At the same time, most of bias mitigation approaches evaluate their performance in English (Zhang et al., 2024b; Peysakhovich & Lerer, 2023; Zhang et al., 2024a; Yu et al., 2024b; Wang et al., 2024b). These approaches fall into two categories: prompt-based techniques and architectural interventions. Architectural methods—such as positional encodings, alternative masking schemes, and calibration mechanisms—address root causes but often require retraining and introduce computational overhead (Zhang et al., 2024b; Wu et al., 2025; Zhao et al., 2021). Prompt-based strategies, including query-aware contextualization and recency prompting, aim to redirect attention dynamically (Peysakhovich & Lerer, 2023; Wang et al., 2024b; Yu et al., 2024b). We focus on prompt-based strategies, using as a starting point the work done by (Zhang et al., 2024a). The authors studied whether a model can improve its performance when given explicit placement of the correct answer. They used two types of instructions — relative and absolute — and found that models lack relative awareness and that implicit information about absolute placement of the correct prompt marginally improves model performance. However, the authors did not study a scenario when the placement prompt is absent at all.
@@
Practical considerations. Position bias affects Chain of Thought Strategies (CoT). CoT struggles with position bias even when reasoning steps are correct, models often fail to retrieve evidence from middle positions (Zhang et al., 2024b). In (Zhang et al., 2024b; Yu et al., 2024b), authors analyzed error propagation in multi-hop reasoning. In RAG systems, one of the mitigating strategies is context ordering  (Wang et al., 2024a; Alessio et al., 2024; Jin et al., 2024) (we discuss these approaches in Appendix A). While conventional approaches often assume a monotonic relationship between document position and attention (e.g., prioritizing the first/last positions), our analysis reveals that position bias may exhibit language-specific patterns and is not always maximized at early tokens. This observation challenges assumptions in methods like Long-Context RAG, which rely on fixed position prioritization, and highlights the need for language-adaptive reordering.
@@
While predictive entropy is widely used to quantify model uncertainty (Huang et al., 2024; Sychev et al., 2025), its relationship with position bias remains unexplored. (Duan et al., 2024) note that uncertainty estimates can be token-biased, but how position bias interacts with uncertainty dynamics is unclear.
@@ 3.2 Context Placement Strategies
Context Volume. Prior research has shown that longer contexts can exacerbate position bias (Baker et al., 2024; Peysakhovich & Lerer, 2023). In our work, we default to five contexts based on preliminary experiments. We vary the number of contexts, setting $N\in\{5,10,15\}$ , and find that while increasing N does degrade performance for some models, the effect is relatively modest. Given limited computational resources, we therefore opted to use five contexts in our main experiments. Preliminary results are provided in Appendix G.
@@ 3.3 Average Predictive Entropy
We adopt the token-wise entropy framework introduced by (Lu et al., 2022), normalized by the total number of tokens, to quantify uncertainty in model responses (Lu et al., 2022; Wang et al., 2025). Let $x$ represent the input prompt and $s=\{z_{1},z_{2},\dots,z_{n}\}$ denote a generated completion sequence of $n$ tokens. For a given large language model (LLM), the conditional probability of generating the $i$-th token $z_{i}$, given the preceding tokens $s_{<i}=\{z_{1},\dots,z_{i-1}\}$ and the prompt $x$, is denoted as $p(z_{i}\mid s_{<i},x)$ for $1\leq i\leq n$. The average predictive entropy1 (denoted as $\textit{PE}_{\text{avg}}$) is defined as:
@@ 4.1 Datasets
Language Source Size
English SQuAD2.0 (Rajpurkar et al., 2018) 150k
Russian MTS-SQuAD (link) 60k
German MLQA (Lewis et al., 2020)  5k
Hindi MLQA (Lewis et al., 2020)  5k
Vietnamese MLQA (Lewis et al., 2020)  5k
@@ 4.2 Models
To investigate whether position bias in large language models (LLMs) arises from model-specific design and training choices or from language-specific characteristics, we evaluate five popular open-source multilingual models: Qwen2.5-7B-Instruct(Qwen et al., 2025), Llama3-8B-Instruct(Touvron et al., 2023), DeepSeek-7B-Chat(DeepSeek-AI, 2024), Gemma-7B-it(Team et al., 2024), and Mistral-7B-Instruct-v0.3(Jiang et al., 2023). All models support English, German, Russian, Vietnamese, and Hindi, yet differ in architecture and training paradigms. This selection enables us to disentangle model driven factors of position bias such as attention mechanisms and training data composition from language-related influences. It is important to note that DeepSeek-7B-Chat performed poorly on Hindi, therefore, we excluded its Hindi results from our analysis.
@@ 5.1 LLM as a Judge
Traditional statistical evaluation metrics of open-ended generations, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and Meteor Universal (Denkowski & Lavie, 2014), are limited in their ability to differentiate between syntactically similar yet semantically distinct texts. Although semantic evaluation methods like BERTScore (Zhang et al., 2019) were developed to overcome these shortcomings, our experimental findings indicate that BERTScore exhibits insufficient discriminative power, frequently failing to capture subtle distinctions between correct and incorrect responses. Consequently, we employ the LLM as a judge framework (Zheng et al., 2023) and select Mistral Large [^2] https://mistral.ai/news/mistral-large-2407 as the evaluator for the following reasons: (1) prior research demonstrates its strong alignment with human judgments and generalizability across diverse tasks (Bavaresco et al., 2024; Kim et al., 2024); (2) it provides a freely accessible API for research, facilitating large-scale evaluation; (3) its architectural design differs from the majority of models used for response generation, thereby reducing potential bias toward self-generated outputs. The evaluator assesses question-answer pairs using a structured prompt that includes the question, ground truth, and model-generated answer. It assigns a label of $1$ for correct answers and 0 for incorrect ones; accuracy is adopted as the primary metric. Further details regarding the prompts and evaluation methodology are provided in Appendix F.
@@ 5.2 Human Evaluation
To validate the reliability of the large language model (LLM) as an evaluative judge, human annotation was conducted on a set of 150 questions in both English and Russian. The responses generated by the Llama model were annotated using the Overlap-3 metric, with domain experts adhering to the same evaluation criteria as the automated judge. Inter-annotator agreement was quantified using Krippendorff’s $\alpha$ (Krippendorff, 2011), yielding a mean $\alpha=0.755$, which indicates a high degree of assessment reliability. Further evaluation of alignment between the automated judge and human annotators is conducted by computing the Pearson correlation coefficient $r$ between the judge’s scores and the majority vote derived from human annotations. A strong mean correlation of $r=0.716$ was observed, indicating substantial agreement. Additional details regarding the annotation procedure are provided in Appendix F.
@@ 6.2 Overall Results Across Languages and Models
Dataset Accuracy
SQuAD2.0 (en) 0.896
MTS-SQuAD (ru) 0.865
MLQA (de) 0.628
MLQA (hi) 0.471
MLQA (vi) 0.475
@@
Model Accuracy
DeepSeek 7B 0.711
Gemma 7B 0.486
Mistral 7B 0.743
Llama3.1 8B 0.690
Qwen2.5 7B 0.754

### v3

@@ Abstract
Large Language Models (LLMs) exhibit position bias systematically underweighting information based on its location in the context but how this bias varies across languages and models remains unclear. We conduct a multilingual study across five typologically diverse languages (English, Russian, German, Hindi, Vietnamese) and five model architectures, analyzing how position bias interacts with prompting strategies and affects output entropy. Our key findings are: (1) Position bias is primarily model-driven but shows language-specific nuances. Notably, Qwen2.5-7B-Instruct, DeepSeek 7B Chat and Mistral 7B consistently favor late positions challenging the common assumption of universal early-token preference. (2) Explicitly instructing the model, in the presence of irrelevant distractors, that “the most relevant context to the query is marked as 1” unexpectedly reduces accuracy across all languages, questioning standard prompt-engineering practices. (3) Accuracy consistently drops most when relevant information appears in the middle of the context, yet this is not reflected in a corresponding increase in output entropy, suggesting the model remains confident even when it fails to use mid-context cues.
Beyond Early-Token Bias: Model-Specific and Language-Specific Position Effects in Multilingual LLMs
Menschikov Mikhail* ITMO Applied AI          Alexander Kharitonov* SberAI          Maiia Kotyga Applied AI          Vadim Porvatov Sber
Anna Zhukovskaya Lomonosov MSU          David Kagramanyan HSE University          Egor Shvetsov Applied AI          Evgeny Burnaev Applied AI AIRI
@@ 1 Introduction
Many recent applications based on large language models (LLMs) require support for long-context processing [^1] Such applications may include Retrieval-Augmented Generation, autonomous agents, scientific research, customer support, among others.. However, developing new training strategies to accommodate longer contexts is insufficient, as new challenges continue to arise. One notable issue is position bias – the systematic neglect of information located at specific positions, typically in the middle of the context (baker2024lost)
While position bias has been well-documented in English-centric studies (zhang2024found; baker2024lost), its manifestation in multilingual contexts and across-various architectures remains underexplored. Furthermore, existing bias mitigation strategies (peysakhovich2023attention; zhao2021calibrate; zhang2024can; guo2024serial) have predominantly been evaluated on English-language datasets. As discussed in Section 2, lexical and morphological variations across languages necessitate careful consideration in multilingual modeling. As a step toward adapting multilingual large language models for real-world long-context applications, this work seeks to address the following research questions:
Q1: Is position bias primarily a model-driven phenomenon, or do language-specific patterns emerge due to lexical, morphological, and syntactic differences (ghosh2024morphology)?
Q2: Do prompt-based strategies, e.g., explicit position guidance (zhang2024can), effectively mitigate bias across languages? Can we focus model attention via prompting, and would it improve model performance?
@@
Contributions. We present a multilingual analysis of position bias across five typologically diverse languages (English, German, Russian, Hindi, and Vietnamese) and five model architectures. Our analysis is based on 2,000 sampled examples per language, where each model was evaluated under nine distinct experimental conditions (3 context positions × 3 scoring strategies), yielding 18,000 model generations and evaluations per language. Across all five languages and five models, this totals 450,000 evaluated question–answer pairs.
Our key contributions are:
  • We show that position bias is predominantly model-driven, yet exhibits significant language-specific variations. For instance, models like Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport), DeepSeek-7B-Chat (deepseek-llm) and Mistral 7B display a strong late-position bias, contradicting prior claims of an inherent early-token preference in LLMs (wu2025emergence; barbero2025llms). In contrast, Llama3.1-8B-Instruct (touvron2023llamaopenefficientfoundation) prioritizes early positions. We speculate these differences stem from variations in training data and model architecture.
  • We find that explicitly instructing models about correct context placement (e.g., “the correct context has label 1”) consistently degrades accuracy across all languages. This result holds when the incorrect context is sampled randomly, a setup that differs from zhang2024can, who report that instructions can mitigate position bias. The key distinction is that zhang2024can use semantically relevant distractors, whereas our distractors are random.
@@
    Retrieval-augmented generation (RAG) systems should account for specific model and language characteristics. Strategies that reorder context based on the assumption that models prioritize recent or initial tokens (see Section 2) may be ineffective or detrimental for models with a late-position bias.
@@
    The observed entropy dynamics complicate uncertainty-based bias mitigation techniques (duan2023shifting) and are crucial for developing effective uncertainty quantification strategies in RAG pipelines.
@@ 2 Related Work
What causes position bias? Prior work identifies multiple contributing factors. (zhang2024found) attribute position bias to U-shaped attention patterns in transformers, which prioritize extremal positions and degrade performance for mid-context evidence. Theoretical and empirical studies further demonstrate that transformers attention is inherently biased toward earlier tokens (wu2025emergence; barbero2025llms). (wu2025emergence) explains position bias due to "causal masking inherently biases attention toward earlier positions, as tokens in deeper layers attend to increasingly contextualized representations of earlier tokens". Our results reveal exceptions where later positions are desired, highlighting the complexity of the problem. Training data biases, such as serial-position effects in corpora, shape how models prioritize sequence positions (wu2025emergence; guo2024serial).
@@
Interplay of Culture, Language, and Model Design. The way we perceive the world is influenced not only by our culture (oyserman2008does) but also by the language we speak (boroditsky2003sex). The latter point is particularly relevant for LLMs, since they are trained on specific languages. Recent studies have shown that multilingual LLMs often initiate their "thinking process" in English, pivoting to the prompt’s original language in the middle layers (zhang2024found; peysakhovich2023attention). These models exhibit lower lexical naturalness in non-English languages, with the naturalness gap being more pronounced for languages structurally distant from English (guo2024large). While the volume of training data plays a crucial role (arnett2024language), linguistic complexity — including lexical and morphological variations across languages — must also be considered (ghosh2024morphology; dang2024morphology; ismayilzada2024evaluating). Additionally, architectural design choices affect languages in different ways; for instance, removing positional encoding from language models would most degrade performance in languages with limited morphological systems (ghosh2024morphology).
@@
At the same time, most of bias mitigation approaches evaluate their performance in English (zhang2024found; peysakhovich2023attention; zhang2024can; yu2024mitigate; wang2024eliminating). These approaches fall into two categories: prompt-based techniques and architectural interventions. Architectural methods—such as positional encodings, alternative masking schemes, and calibration mechanisms—address root causes but often require retraining and introduce computational overhead (zhang2024found; wu2025emergence; zhao2021calibrate). Prompt-based strategies, including query-aware contextualization and recency prompting, aim to redirect attention dynamically (peysakhovich2023attention; wang2024eliminating; yu2024mitigate). We focus on prompt-based strategies, using as a starting point the work done by (zhang2024can). The authors studied whether a model can improve its performance when given explicit placement of the correct answer. They used two types of instructions — relative and absolute — and found that models lack relative awareness and that implicit information about absolute placement of the correct prompt marginally improves model performance. The main difference with this approach and ours is that authors in  (zhang2024can) use relevant distractors, while in our work we use random distractors.
@@
Practical considerations. Position bias affects Chain of Thought Strategies (CoT). CoT struggles with position bias even when reasoning steps are correct, models often fail to retrieve evidence from middle positions (zhang2024found). In (zhang2024found; yu2024mitigate), authors analyzed error propagation in multi-hop reasoning. In RAG systems, one of the mitigating strategies is context ordering  (wang2024rear; alessio2024improving; jin2024long) (we discuss these approaches in Appendix A). While conventional approaches often assume a monotonic relationship between document position and attention (e.g., prioritizing the first/last positions), our analysis reveals that position bias may exhibit language-specific patterns and is not always maximized at early tokens. This observation challenges assumptions in methods like Long-Context RAG, which rely on fixed position prioritization, and highlights the need for language-adaptive reordering.
@@
While predictive entropy is widely used to quantify model uncertainty (huang2024survey; sychev2025llm), its relationship with position bias remains unexplored. (duan2023shifting) note that uncertainty estimates can be token-biased, but how position bias interacts with uncertainty dynamics is unclear.
@@ 3.2 Context Placement Strategies
Context Volume. Prior research has shown that longer contexts can exacerbate position bias (baker2024lost; peysakhovich2023attention). In our work, we default to five contexts based on preliminary experiments. We vary the number of contexts, setting $N\in\{5,10,15\}$ , and find that while increasing N does degrade performance for some models, the effect is relatively modest. Given limited computational resources, we therefore opted to use five contexts in our main experiments. Preliminary results are provided in Appendix G.
@@ 3.3 Average Predictive Entropy
We adopt the token-wise entropy framework introduced by (lu2021fantastically), normalized by the total number of tokens, to quantify uncertainty in model responses (lu2021fantastically; wang2025word). Let $x$ represent the input prompt and $s=\{z_{1},z_{2},\dots,z_{n}\}$ denote a generated completion sequence of $n$ tokens. For a given large language model (LLM), the conditional probability of generating the $i$-th token $z_{i}$, given the preceding tokens $s_{<i}=\{z_{1},\dots,z_{i-1}\}$ and the prompt $x$, is denoted as $p(z_{i}\mid s_{<i},x)$ for $1\leq i\leq n$. The average predictive entropy1 (denoted as $\textit{PE}_{\text{avg}}$) is defined as:
@@ 4.1 Datasets
Language Source Size
English SQuAD2.0 (rajpurkar2018know) 150k
Russian MTS-SQuAD (link) 60k
German MLQA (lewis2020mlqa)  5k
Hindi MLQA (lewis2020mlqa)  5k
Vietnamese MLQA (lewis2020mlqa)  5k
@@ 4.2 Models
To investigate whether position bias in large language models (LLMs) arises from model-specific design and training choices or from language-specific characteristics, we evaluate five popular open-source multilingual models: Qwen2.5-7B-Instruct(qwen2025qwen25technicalreport), Llama3-8B-Instruct(touvron2023llamaopenefficientfoundation), DeepSeek-7B-Chat(deepseek-llm), Gemma-7B-it(gemmateam2024gemmaopenmodelsbased), and Mistral-7B-Instruct-v0.3(jiang2023mistral7b). All models support English, German, Russian, Vietnamese, and Hindi, yet differ in architecture and training paradigms. This selection enables us to disentangle model driven factors of position bias such as attention mechanisms and training data composition from language-related influences. It is important to note that DeepSeek-7B-Chat performed poorly on Hindi, therefore, we excluded its Hindi results from our analysis.
@@ 5.1 LLM as a Judge
Traditional statistical evaluation metrics of open-ended generations, such as BLEU (papineni2002bleu), ROUGE (lin2004rouge), and Meteor Universal (denkowski2014meteor), are limited in their ability to differentiate between syntactically similar yet semantically distinct texts. Although semantic evaluation methods like BERTScore (zhang2019bertscore) were developed to overcome these shortcomings, our experimental findings indicate that BERTScore exhibits insufficient discriminative power, frequently failing to capture subtle distinctions between correct and incorrect responses. Consequently, we employ the LLM as a judge framework (zheng2023judging) and select Mistral Large [^2] https://mistral.ai/news/mistral-large-2407 as the evaluator for the following reasons: (1) prior research demonstrates its strong alignment with human judgments and generalizability across diverse tasks (bavaresco2406llms; kim2024prometheus); (2) it provides a freely accessible API for research, facilitating large-scale evaluation; (3) its architectural design differs from the majority of models used for response generation, thereby reducing potential bias toward self-generated outputs. The evaluator assesses question-answer pairs using a structured prompt that includes the question, ground truth, and model-generated answer. It assigns a label of $1$ for correct answers and 0 for incorrect ones; accuracy is adopted as the primary metric. Further details regarding the prompts and evaluation methodology are provided in Appendix F.
@@ 5.2 Human Evaluation
To validate the reliability of the large language model (LLM) as an evaluative judge, human annotation was conducted on a set of 150 questions in both English and Russian. The responses generated by the Llama model were annotated using the Overlap-3 metric, with domain experts adhering to the same evaluation criteria as the automated judge. Inter-annotator agreement was quantified using Krippendorff’s $\alpha$ (krippendorff2011computing), yielding a mean $\alpha=0.755$, which indicates a high degree of assessment reliability. Further evaluation of alignment between the automated judge and human annotators is conducted by computing the Pearson correlation coefficient $r$ between the judge’s scores and the majority vote derived from human annotations. A strong mean correlation of $r=0.716$ was observed, indicating substantial agreement. Additional details regarding the annotation procedure are provided in Appendix F.
@@ 6.2 Overall Results Across Languages and Models
Dataset Acc
SQuAD2.0 (en) 89.6
MTS-SQuAD (ru) 86.5
MLQA (de) 62.8
MLQA (hi) 47.1
MLQA (vi) 47.5
@@
Model Acc
DeepSeek 7B 71.1
Gemma 7B 48.6
Mistral 7B 74.3
Llama3.1 8B 69.0
Qwen2.5 7B 75.4